THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 103 — 1 February 2010

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues,

It is my pleasure to present you the 103^{rd} issue of the Magellanic Clouds Newsletter. Star clusters, star forming regions, X-ray binaries and supernova remnants feature prominently, but the other papers are well worth reading too. A great start to 2010!

There is no reason why we should not have a cover picture. So: if you have a beautiful image, diagram, or spectrum related to the Magellanic Clouds, why not send it to mcnews@astro.keele.ac.uk and ask it to be considered to be on the cover of one of the next issues? It could advertise your paper, work in progress, or just be something to please the eye. Postscript is the (highly) preferred format.

Figure 1: Spitzer Space Telescope far-infrared spectrum of LMC supernova remnant N 49 (AJ, 139, 68 (2010)).

The next issue is planned to be distributed on the 1^{st} of April 2010.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

VLT/NACO near-infrared imaging and spectroscopy of N 88A in the SMC

G. Testor^{1}, J.L. Lemaire^{1,2}, M. Heydari-Malayeri^{1}, L.E. Kristensen^{3}, S. Diana^{2} and D. Field^{4}

1LERMA, UMR 8112 du CNRS, Observatoire de Paris, 92195 Meudon, France 2Université de Cergy-Pontoise, 95031 Cergy Cedex, France 3Leiden Observatory, Leiden University, Niels Bohrweg 2, 2333 CA Leiden, The Netherlands 4Department of Physics and Astronomy, Århus University, 8000 Århus C, Denmark

We present near-infrared imaging and spectroscopic high spatial resolution observations of the SMC region N 88 containing the bright, excited, extincted and compact H ii region N 88A of the size of about ∼ 1 pc. To investigate its stellar content and reddening, N 88 was observed using spectroscopy and imagery in the JHK_{s}- and L′-band at a spatial resolution of ∼ 0.1–0.3′′, using the VLT UT4 equipped with the NAOS adaptive optics system. In an attempt to establish if the origin of the infra-red (IR) excess is due to bright nebulosity, circumstellar material and/or local dust, we used K_{s} vs. J–K colour–magnitude (CM) and JHK colour–colour (CC) diagrams, as well as L′ imagery. Our IR-data reveal in the N 88 area an IR-excess fraction of ≥ 30 per cent of the detected stars as well as an unprecedently detailed morphology of N 88A. It consists of an embedded cluster of ∼ 3.5′′ (∼ 1 pc) in diameter of at least thirteen resolved stars superposed with an unusual bright continuum centered on a very bright star. The four brightest stars in this cluster lie red-ward of H − K ≥ 0.45 mag and could be classified as young stellar object (YSO) candidates. Four other probable YSO candidates are also detected in N88 along a north-south bow-shaped thin H_{2} filament at ∼ 7′′ east of the young central bright star. This star, which we assume to be the main exciting source, could also be complex. At 0.2′′ east of this star, a heavily embedded core is detected in the L′-band. This core with L′ ∼ 14 mag and L′ − K ≥ 4.5 mag could be a massive class I protostar candidate. The 2.12 μm H_{2} image of N88A resembles a shell of a diameter of ∼ 3′′ (∼ 0.9 pc) centered on the bright star. This shell consists of three bright components, of which the brightest one superposes the ionization front. The line ratios of H_{2} 2–1 S(1) and 1–0 S(0) relative to 1–0 S(1), as well as the presence of high v lines, are indicative of photodissociation regions, rather than shocks.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:0911.5490

Optical spectroscopy of 20 Be/X-ray Binaries in the Small Magellanic Cloud

Vallia Antoniou^{1,2}, Despina Hatzidimitriou^{1}, Andreas Zezas^{1,2,3} and Pablo Reig^{1,3}

1Physics Department, University of Crete, P.O. Box 2208, GR-710 03, Heraklion, Crete, Greece 2Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA; vantoniou@head.cfa.harvard.edu 3IESL, Foundation for Research and Technology, 71110 Heraklion, Crete, Greece

We present a large sample (20 in total) of optical spectra of Small Magellanic Cloud (SMC) High-Mass X-ray Binaries obtained with the 2dF spectrograph at the Anglo-Australian Telescope. All of these sources are found to be Be/X-ray binaries (Be-XRBs), while for 5 sources we present original classifications. Several statistical tests on this expanded sample support previous findings for similar spectral-type distributions of Be-XRBs and Be field stars in the SMC, and of Be-XRBs in the Large Magellanic Cloud and the Milky Way, although this could be the result of small samples. On the other hand, we find that Be-XRBs follow a different distribution than Be stars in the Galaxy, also in agreement with previous studies. In addition, we find similar Be spectral type distributions between the Magellanic Clouds samples. These results reinforce the relation between the orbital period and the equivalent width of the Hα line that holds for Be-XRBs. SMC Be stars have larger Hα equivalent widths when compared to Be-XRBs, supporting the notion of circumstellar disk truncation by the compact object.

Published in The Astrophysical Journal, 707, 1080 (2009) Available from arXiv:0912.1876

The star formation history of the SMC star cluster NGC 419

Stefano Rubele ^{1,2}, Leandro Kerber ^{3} and Léo Girardi^{1}

1Oss. Ast. Padova, Italy 2UESC, Ilhéus, Brazil 3Dip. Astro. Padova, Italy

The rich SMC star cluster NGC 419 has recently been found to present both a broad main sequence turn-off and a dual red clump of giants, in the sharp colour–magnitude diagrams (CMD) derived from the High Resolution Channel of the Advanced Camera for Surveys on board the Hubble Space Telescope. In this work, we apply to the NGC 419 data the classical method of star formation history (SFH) recovery via CMD reconstruction, deriving for the first time this function for a star cluster with multiple turn-offs. The values for the cluster metallicity, reddening, distance and binary fraction, were varied within the limits allowed by present observations. The global best-fitting solution is an excellent fit to the data, reproducing all the CMD features with striking accuracy. The corresponding star formation rate is provided together with estimates of its random and systematic errors. Star formation is found to last for at least 700 Myr, and to have a marked peak at the middle of this interval, for an age of 1.5 Gyr. Our findings argue in favour of multiple star formation episodes (or continued star formation) being at the origin of the multiple main sequence turn-offs in Magellanic Cloud clusters with ages around 1 Gyr. It remains to be tested whether alternative hypotheses, such as a main sequence spread caused by rotation, could produce similarly good fits to the data.

Accepted for publication in MNRAS Available from arXiv:0912.2638

Hierarchical structures in the Large and Small Magellanic Clouds

Charles Bonatto^{1} and Eduardo Bica^{1}

1Departamento de Astronomia, Universidade Federal do Rio Grande do Sul, Av. Bento Gonçalves 9500, Porto Alegre 91501-970, RS, Brazil

We investigate the degree of spatial correlation among extended structures in the LMC and SMC. To this purpose we work with sub-samples characterised by different properties such as age and size, taken from the updated catalogue of Bica et al. or gathered in the present work. The structures are classified as star clusters or non-clusters (basically, nebular complexes and their stellar associations). The radius distribution functions follow power-laws (dN/dR ∝ R^{−α}) with slopes and maximum radius (R_{max}) that depend on object class (and age). Non-clusters are characterised by α ≈ 1.9 and R_{max}< ∼472 pc, while young clusters (age < ∼10 Myr) have α ≈ 3.6 and R_{max}< ∼15 pc, and old ones (age > ∼600 Myr) have α ≈ 2.5 and R_{max}< ∼40 pc. Young clusters present a high degree of spatial self-correlation and, especially, correlate with star-forming structures, which does not occur with the old ones. This is consistent with the old clusters having been heavily mixed up, since their ages correspond to several LMC and SMC crossing times. On the other hand, with ages corresponding to fractions of the respective crossing times, the young clusters still trace most of their birthplace structural pattern. Also, small clusters (R < 10 pc), as well as small non-clusters (R < 100 pc), are spatially self-correlated, while their large counterparts of both classes are not. The above results are consistent with a hierarchical star-formation scenario for the LMC and SMC.

Accepted for publication in MNRAS Available from arXiv:0912.1745

Pre-main sequence stars in the stellar association N 11 in the Large Magellanic Cloud

A. Vallenari^{1}, E. Chiosi^{1} and R. Sordo^{1}

1INAF, OAPD, Italy

Magellanic Clouds are of extreme importance to the study of the star formation process in low metallicity environments. In this paper we report on the discovery of pre-main sequence candidates and young embedded stellar objects in N 11 located in the Large Magellanic Cloud to cast light on the star formation scenario. We would like to remind that this comparison is complicated by the presence of a large age dispersion detected in the fields. Deep archive HST/ACS photometry is used to derive color–magnitude diagrams of the associations in N 11 and of the foreground field population. These data are complemented by archive IR Spitzer data which allow the detection of young embedded stellar objects. The spatial distribution of the pre-main sequence candidates and young embedded stellar objects is compared with literature data observed at different wavelengths, such as Hα and CO maps, and with the distribution of OB and Herbig Ae/Be stars. The degree of clustering is derived using the Minimal Spanning Tree method and the two point correlation function to get insights about the formation process. A large population of pre-main sequence candidates is found in N 11. Their masses are in the range of 1.3–2 M_{⊙} for ages from 2 to 10 Myr. Young embedded stellar objects having ages of 0.1–1 Myr are found to be intermixed with the candidate pre-main sequence stars. The spatial distribution of the stars shows that this region is the product of clustered star formation. No significant difference is found in the clustering degree of young blue main sequence stars and faint pre-main sequence candidates, suggesting that they might be part of the same formation process. The data suggest that the star formation in the region is a long-lasting process where stars from 0.1 to 10 Myr are widely distributed.

Accepted for publication in A&A Available from arXiv:0912.0176

Be/X-ray binary SXP 6.85 undergoes large Type II outburst in the Small Magellanic Cloud

L.J. Townsend^{1}, M.J. Coe^{1}, V.A. McBride^{1}, A.J. Bird^{1}, M.P.E. Schurch^{2}, R.H.D. Corbet^{3}, F. Haberl^{4}, J.L. Galache^{5} and A. Udalski^{6}

1School of Physics and Astronomy, University of Southampton, Highfield, Southampton, SO17 1BJ, United Kingdom 2Department of Astronomy, University of Cape Town, Private Bag X3, Rondebosch 7701, South Africa 3University of Maryland Baltimore County, X-ray Astrophysics Laboratory, Mail Code 662, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA 4Max-Planck-Institut für extraterrestrische Physik, Gießenbachstraße, 85748, Germany 5Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, USA 6Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warsaw, Poland

The Small Magellanic Cloud (SMC) Be/X-ray binary pulsar SXP 6.85 = XTE J0103−728 underwent a large Type II outburst beginning on 2008 August 10. The source was consistently seen for the following 20 weeks (MJD = 54688–54830). We present X-ray timing and spectroscopic analysis of the source as part of our ongoing Rossi X-ray Timing Explorer (RXTE) monitoring campaign and INTEGRAL key programme monitoring the SMC and 47 Tuc. A comparison with the Optical Gravitational Lensing Experiment (OGLE) III light curve of the Be counterpart shows the X-ray outbursts from this source coincide with times of optical maximum. We attribute this to the circumstellar disk increasing in size, causing mass accretion onto the neutron star. Ground based IR photometry and Hα spectroscopy obtained during the outburst are used as a measure of the size of the circumstellar disk and lend support to this picture. In addition, folded RXTE light curves seem to indicate complex changes in the geometry of the accretion regions on the surface of the neutron star, which may be indicative of an inhomogeneous density distribution in the circumstellar material causing a variable accretion rate onto the neutron star. Finally, the assumed inclination of the system and Hα equivalent width measurements are used to make a simplistic estimate of the size of the circumstellar disk.

Accepted for publication in MNRAS Available from arXiv:0912.2940

The Magellanic Bridge: evidence for a population of X-ray binaries

V.A. McBride^{1}, A.J. Bird^{1}, M.J. Coe^{1}, L.J. Townsend^{1}, R.H.D. Corbet^{2,3} and F. Haberl^{4}

1University of Southampton, UK 2University of Maryland, USA 3CRESST, NASA Goddard Space Flight Center, USA 4Max-Planck-Institut für extraterrestrische Physik, Germany

INTEGRAL observations of the Small Magellanic Cloud region have resulted in the serendipitous detection of two transient hard X-ray sources in the Magellanic Bridge. In this paper we present the timing and spectral characteristics of these sources across the 2–100 keV energy range, which, in conjunction with their optical counterparts, demonstrate that they are high mass X-ray binaries in the Magellanic Bridge. Together with one previously known high mass X-ray binary system, and three candidates, these sources represent an emerging population of X-ray binaries in the Bridge, probably initiated by tidally induced star formation as a result of the gravitational interaction between the Large and Small Magellanic Clouds.

Accepted for publication in MNRAS Available from arXiv:0912.2951

Ca ii triplet spectroscopy of Small Magellanic Cloud red giants. II. Abundances for a sample of field stars.

M.C. Parisi^{1}, D. Geisler^{2}, A.J. Grocholski^{3}, J.J. Clariá^{1} and A. Sarajedini^{4}

1Observatorio Astronómico, Universidad Nacional de Córdoba, Argentina 2Departamento de Astronomía, Universidad de Concepción, Chile 3Space Telescope Science Institute, USA 4Department of Astronomy, University of Florida, USA

We have obtained metallicities of ∼ 360 red giant stars distributed in 15 Small Magellanic Cloud (SMC) fields from near-infrared spectra covering the Ca ii triplet lines using the VLT + FORS2. The errors of the derived [Fe/H] values range from 0.09 to 0.35 dex per star, with a mean of 0.17 dex. The metallicity distribution of the whole sample shows a mean value of [Fe/H] = −1.00 ± 0.02, with a dispersion of 0.32 ± 0.01, in agreement with global mean [Fe/H] values found in previous studies. We find no evidence of a metallicity gradient in the SMC. In fact, on analysing the metallicity distribution of each field, we derived mean values of [Fe/H] = −0.99 ± 0.08 and [Fe/H] = −1.02 ± 0.07 for fields located closer and farther than 4^{∘} from the center of the galaxy, respectively. In addition, there is a clear tendency for the field stars to be more metal-poor than the corresponding cluster they surround, independent of their positions in the galaxy and of the clusters’ age. We argue that this most likely stems from the field stars being somewhat older and therefore somewhat more metal-poor than most of our clusters.

Accepted for publication in Astronomical Journal

Multi-frequency radio measurements of SN 1987A over 22 Years

G. Zanardo^{1}, L. Staveley-Smith^{1}, Lewis Ball^{2}, B.M. Gaensler^{3}, M.J. Kesteven^{2}, R.N. Manchester^{2}, C.-Y. Ng^{3}, A.K. Tzioumis^{2} and T.M. Potter^{1}

1International Centre for Radio Astronomy Research (ICRAR), School of Physics M013, The University of Western Australia, Crawley, WA 6009, Australia 2CSIRO Australia Telescope National Facility, PO Box 76, Epping, NSW 1710, Australia 3Sydney Institute for Astronomy (SIfA), School of Physics, The University of Sydney, NSW 2006, Australia

We present extensive observations of the radio emission from the remnant of SN 1987A made with the Australia Telescope Compact Array (ATCA), since the first detection of the remnant in 1990. The radio emission has evolved in time providing unique information on the interaction of the supernova shock with the circumstellar medium. We particularly focus on the monitoring observations at 1.4, 2.4, 4.8 and 8.6 GHz, which have been made at intervals of 4–6 weeks. The flux density data show that the remnant brightness is now increasing exponentially, while the radio spectrum is flattening. The current spectral index value of −0.68 represents an 18 ± 3% increase over the last 8 years. The exponential trend in the flux is also found in the ATCA imaging observations at 9 GHz, which have been made since 1992, approximately twice a year, as well as in the 843 MHz data set from the Molonglo Observatory Synthesis Telescope from 1987 to March 2007. Comparisons with data at different wavelengths (X-ray, Hα) are made. The rich data set that has been assembled in the last 22 years forms a basis for a better understanding of the evolution of the supernova remnant.

Accepted for publication in The Astrophysical Journal Available from arXiv:0912.4979

Toward the general RGB slope–metallicity-age calibration: I Metallicities, ages and kinematics for eight LMC clusters

Saurabh Sharma^{1}, Jura Bosrissova^{1}, Radostin Kurtev^{1}, V.D. Ivanov^{2} and D. Geisler^{3}

1Departamento de Física y Astronomía, Facultad de Ciencias, Universidad de Valparaiso, Ave. Gran Bretaña 1111, Playa Ancha, Casilla 53, Valparaiso, Chile 2European Southern Observatory, Ave. Alonso de Cordoba 3107, Casilla 19, Santiago 19001, Chile 3Departamento de Astronomía, Universidad de Concepción, Casilla 160-C, Concepción, Chile

In this paper, we discuss the properties of CMDs, age, metallicity and radial velocities of eight massive LMC clusters using data taken from FORS2 multiobject spectrograph at the 8.2-meter VLT/UT1. The strong near-infrared Ca ii triplet (CaT) lines of RGB stars obtained from the high S/N spectra are used to determine the metallicity and radial velocity of cluster members. We report for the first time spectroscopically determined metallicity values for four clusters based on the mean [Fe/H] value of ∼ 10 cluster members each. We found two concentrations in the distribution of ages of the target clusters. Six have ages between 0.8–2.2 Gyr and the other two, NGC 1754 and NGC 1786, are very old. The metallicity of the six intermediate age clusters, with a mean age of 1.5 Gyr, is −0.49 with a scatter of only 0.04. This tight distribution suggests that a close encounter between the LMC and SMC may have caused not only the restart of cluster formation in the LMC but the generation of the central bar. The metallicity for the two old clusters is similar to that of the other old, metal-poor LMC clusters. We find that the LMC cluster system exhibits disk-like rotation with no clusters appearing to have halo kinematics and there is no evidence of a metallicity gradient in the LMC, in contrast with the stellar population of the MW and M 33, where the metallicity decreases as galactocentric distance increases. The LMC’s stellar bar may be the factor responsible for the dilution of any kind of gradient in the LMC.

Accepted for publication in Astronomical Journal Available from arXiv:0912.5208

The Optical Gravitational Lensing Experiment. The OGLE-III Catalog of Variable Stars. V. R Coronae Borealis stars in the Large Magellanic

Cloud

I. Soszyński^{1}, A. Udalski^{1}, M.K. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, Ł. Wyrzykowski^{3}, O. Szewczyk^{2}, K. Ulaczyk^{1} and R. Poleski^{1}

1Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2Universidad de Concepción, Departamento de Física, Casilla 160-C, Concepción, Chile 3Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK

The fifth part of the OGLE-III Catalog of Variable Stars presents 23 R CrB (RCB) stars in the Large Magellanic Cloud (LMC). 17 of these objects have been spectroscopically confirmed by previous studies, while 6 stars are new candidates for RCB variables. We publish the VI multi-epoch OGLE photometry for all objects. We use the sample of carbon-rich long-period variables released in the previous part of this catalog to select objects with severe drops in luminosity, i.e. with the DY Per-like light curves. DY Per stars are often related to R CrB variables. We detect at least 600 candidates for DY Per stars, mostly among dust enshrouded giants. We notice that our candidate DY Per stars form a continuity with other carbon-rich long-period variables, so it seems that DY Per stars do not constitute a separate group of variable stars.

Published in Acta Astronomica Available from arXiv:0912.2097 and from http://acta.astrouw.edu.pl/Vol59/n4/pdf/pap 59 4 1.pdf

Assessment of stellar stratification in three young star clusters in the Large Magellanic Cloud

Dimitrios Gouliermis^{1}, Dougal Mackey^{2}, Yu Xin^{3} and Boyke Rochau^{1}

1Max Planck Institute for Astronomy, Königstuhl 17, 69117 Heidelberg, Germany 2Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK 3Argelander-Institut für Astronomie, Rheinische Friedrich-Wilhelms-Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany

We present a comprehensive study of stellar stratification in young star clusters in the Large Magellanic Cloud (LMC). We apply our recently developed effective radius method for the assessment of stellar stratification on imaging data obtained with the Advanced Camera for Surveys of three young LMC clusters to characterize the phenomenon and develop a comparative scheme for its assessment in such clusters. The clusters of our sample, NGC 1983, NGC 2002 and NGC 2010, are selected on the basis of their youthfulness, and their variety in appearance, structure, stellar content, and surrounding stellar ambient. Our photometry is complete for magnitudes down to m_{814} ≃ 23 mag, allowing the calculation of the structural parameters of the clusters, the estimation of their ages and the determination of their stellar content. Our study shows that each cluster in our sample demonstrates stellar stratification in a quite different manner and at different degree from the others. Specifically, NGC 1983 shows to be partially segregated with the effective radius increasing with fainter magnitudes only for the faintest stars of the cluster. Our method on NGC 2002 provides evidence of strong stellar stratification for both bright and faint stars; the cluster demonstrates the phenomenon with the highest degree in the sample. Finally, NGC 2010 is not segregated, as its bright stellar content is not centrally concentrated, the relation of effective radius to magnitude for stars of intermediate brightness is rather flat, and we find no evidence of stratification for its faintest stars. For the parameterization of the phenomenon of stellar stratification and its quantitative comparison among these clusters, we propose the slope derived from the change in the effective radius over the corresponding magnitude range as indicative parameter of the degree of stratification in the clusters. A positive value of this slope indicates mass segregation in the cluster, while a negative or zero value signifies the lack of the phenomenon.

Published in Astrophysical Journal, 709, 263 (2010) Available from arXiv:1001.0135 and from http://rapidshare.com/files/328406139/Gouliermis 2010.ApJ.709.pdf

A deep Chandra observation of the oxygen-rich supernova remnant 0540−69.3 in the Large Magellanic Cloud

Sangwook Park^{1}, John P. Hughes^{2}, Patrick O. Slane^{3}, Koji Mori^{4} and David N. Burrows^{1}

1Penn State, USA 2Rutgers, USA 3CfA, USA 4Miyazaki, Japan

Using our deep ∼120 ks Chandra observation, we report on the results from our spatially-resolved X-ray spectral analysis of the “oxygen-rich” supernova remnant (SNR) 0540−69.3 in the Large Magellanic Cloud. We conclusively establish the nonthermal nature of the “arcs” in the east and west boundaries of the SNR, which confirms the cosmicray electron acceleration in the supernova shock (B ∼ 20–140 μG). We report tentative evidence for Fe overabundance in the southern region close to the outer boundary of the SNR. While such a detection would be intriguing, the existence of Fe ejecta is not conclusive with the current data because of poor photon statistics and limited plasma models. If it is verified using deeper X-ray observations and improved plasma models, the presence of Fe ejecta, which was produced in the core of the supernova, near the SNR’s outer boundary would provide an intriguing opportunity to study the explosive nucleosynthesis and the ejecta mixing in this young core-collapse SNR. There is no evidence of X-ray counterparts for the optical O-rich ejecta in the central regions of the SNR.

Accepted for publication in ApJ Available from arXiv:0912.5177

Masers associated with high-mass star formation regions in the Large Magellanic Cloud

S.P. Ellingsen^{1}, S.L. Breen^{1,2}, J.L. Caswell^{2}, L.J. Quinn^{3} and G.A. Fuller^{3}

1School of Mathematics and Physics, University of Tasmania, Australia 2Australia Telescope National Facility, CSIRO, Australia 3Jodrell Bank Centre for Astrophysics, University of Manchester, UK

We report the results of a sensitive search for 12.2-GHz methanol maser emission towards a sample of eight highmass star formation regions in the Large Magellanic Clouds which have been detected in other maser transitions. We detected one source towards the star formation region N 105a. This is the first detection of a 12.2-GHz methanol maser outside our Galaxy. We also made near-contemporaneous observations of the 6.7-GHz methanol and 22-GHz water masers towards these sources, resulting in the detection of water maser emission in six new sources, including one associated with the strongest 6.7-GHz maser in the Magellanic Clouds IRAS 05011−6815. The majority of the maser sources are closely associated with objects identified as likely Young Stellar Objects (YSO) on the basis of Spitzer Space Telescope observations. We find that the YSOs associated with masers tend to be more luminous and have redder infrared colours than the sample as a whole. SED modeling of the YSOs shows that the masers are associated with sources of higher central mass, total luminosity and ambient density than the majority of YSOs in the LMC. This is consistent with the well-established relationship between luminous methanol and water masers and young, high-mass objects observed in the Galaxy.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1001.2169

AKARI near- to mid-infrared imaging and spectroscopic observations of the Small Magellanic Cloud. I. Bright point source list

Yoshifusa Ita^{1}, T. Onaka^{2}, T. Tanabé^{3}, N. Matsunaga^{3}, M. Matsuura^{4}, I. Yamamura^{5}, Y. Nakada^{3}, H. Izumiura^{6}, T. Ueta^{7}, H. Mito^{8}, H. Fukushi^{3} and D. Kato^{2}

1NAOJ, Japan 2Department of Astronomy, University of Tokyo, Japan 3Institute of Astronomy, University of Tokyo, Japan 4UCL, UK 5ISAS/JAXA, Japan 6OAO/NAOJ, Japan 7University of Denver, USA 8Kiso Observatory, University of Tokyo, Japan

We carried out a near- to mid-infrared imaging and spectroscopic observations of the patchy areas in the Small Magellanic Cloud using the Infrared Camera on board AKARI. Two 100 arcmin^{2} areas were imaged in 3.2, 4.1, 7, 11, 15, and 24 μm and also spectroscopically observed in the wavelength range continuously from 2.5 to 13.4 μm. The spectral resolving power (λ/Δλ) is about 20, 50, and 50 at 3.5, 6.6 and 10.6 μm, respectively. Other than the two 100 arcmin^{2} areas, some patchy areas were imaged and/or spectroscopically observed as well. In this paper, we overview the observations and present a list of near- to mid-infrared photometric results, which lists ∼ 12, 000 near-infrared and ∼ 1, 800 mid-infrared bright point sources detected in the observed areas. The 10-σ limits are 16.50, 16.12, 13.28, 11.26, 9.62, and 8.76 in Vega magnitudes at 3.2, 4.1, 7, 11, 15, and 24 μm bands, respectively.

Accepted for publication in PASJ Available from arXiv:1001.1791 and from http://www-irc.mtk.nao.ac.jp/%7Eyita/smc20100112.pdf

The Onfp Class in the Magellanic Clouds

Nolan R. Walborn^{1}, Ian D. Howarth^{2}, Christopher J. Evans^{3}, Paul A. Crowther^{4}, Anthony F. J. Moffat^{5}, Nicole St-Louis^{5}, Cecilia Fariña^{6}, Guillermo L. Bosch^{6}, Nidia I. Morrell^{7}, Rodolfo H. Barbá^{8,9} and Jacco Th. van Loon^{10}

1Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 3UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh EH9 3HJ, UK 4Department of Physics and Astronomy, University of Sheffield, Hounsfield Road, Sheffield S3 7RH, UK 5Département de Physique, Université de Montreal, C.P. 6128, Succ. Centre-Ville, Montreal, QC H3C 3J7, Canada 6Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, 1900 La Plata, Argentina and IALP-CONICET, Argentina 7Las Campanas Observatory, Observatories of the Carnegie Institution of Washington, Casilla 601, La Serena, Chile 8Departamento de Física, Universidad de La Serena, Cisternas 1200 Norte, La Serena, Chile 9Instituto de Ciencias Astronómicas de la Tierra y del Espacio (ICATE–CONICET), Avenida España 1512 Sur, J5402DSP, San Juan, Argentina 10Astrophysics Group, School of Physical & Geographical Sciences, Keele University, Staffordshire ST5 5BG, UK

The Onfp class of rotationally broadened, hot spectra was defined some time ago in the Galaxy, where its membership to date numbers only eight. The principal defining characteristic is a broad, centrally reversed He ii λ4686 emission profile; other emission and absorption lines are also rotationally broadened. Recent surveys in the Magellanic Clouds (MCs) have brought the class membership there, including some related spectra, to 28. We present a survey of the spectral morphology and rotational velocities, as a first step toward elucidating the nature of this class. Evolved, rapidly rotating hot stars are not expected theoretically, because the stellar winds should brake the rotation. Luminosity classification of these spectra is not possible, because the principal criterion (He ii λ4686) is peculiar; however, the MCs provide reliable absolute magnitudes, which show that they span the entire range from dwarfs to supergiants. The Onfp line-broadening distribution is distinct and shifted toward larger values from those of normal O dwarfs and supergiants with >99.99% confidence. All cases with multiple observations show line-profile variations, which even remove some objects from the class temporarily. Some of them are spectroscopic binaries; it is possible that the peculiar profiles may have multiple causes among different objects. The origin and future of these stars are intriguing; for instance, they could be stellar mergers and/or γ-ray-burst progenitors.

Accepted for publication in The Astronomical Journal Available from arXiv:1001.4032

Type II Cepheids as extragalactic distance candles

Daniel J. Majaess^{1,2}, David G. Turner^{1,2} and David J. Lane^{1,2}

1Saint Mary’s University, Halifax, Nova Scotia, Canada 2The Abbey Ridge Observatory, Stillwater Lake, Nova Scotia, Canada

Extragalactic type II Cepheids are tentatively identified in photometric surveys of IC 1613, M 33, M 101, M 106, M 31, NGC 4603, and the SMC. Preliminary results suggest that type II Cepheids may play an important role as standard candles, in constraining the effects of metallicity on Cepheid parameters, and in mapping extinction.

Published in Acta Astronomica Available from arXiv:0909.0181 and from http://acta.astrouw.edu.pl/Vol59/n4/a 59 4 5.html

Interstellar Ti ii in the Milky Way and Magellanic Clouds

Daniel E. Welty^{1} and Paul A. Crowther^{2}

1University of Illinois at Urbana/Champaign, USA 2University of Sheffield, UK

We discuss several sets of Ti ii absorption-line data, which probe a variety of interstellar environments in our Galaxy and in the Magellanic Clouds. Comparisons of high-resolution (FWHM ∼ 1.3–1.5 km s^{−1}) Ti ii spectra of Galactic targets with corresponding high-resolution spectra of Na i, K i, and Ca ii reveal both similarities and differences in the detailed structure of the absorption-line profiles — reflecting component-to-component differences in the ionization and depletion behaviour of those species. Moderate-resolution (FWHM ∼ 3.4–4.5 km s^{−1}) spectra of more heavily reddened Galactic stars provide more extensive information on the titanium depletion in colder, denser clouds — where more than 99.9 per cent of the Ti may be in the dust phase. Moderate-resolution (FWHM ∼ 4.5–8.7 km s^{−1}) spectra of stars in the Magellanic Clouds suggest that the titanium depletion is generally much less severe in the LMC and SMC than in our Galaxy [for a given N(H_{tot}), E(B −V), or molecular fraction f(H_{2})] — providing additional evidence for differences in depletion patterns in those two lower-metallicity galaxies. We briefly discuss possible implications of these results for the interpretation of gas-phase abundances in QSO absorption-line systems and of variations in the D/H ratio in the local Galactic ISM.

Accepted for publication in MNRAS Available from arXiv:1001.3687

A Spitzer Space Telescope far-infrared spectral atlas of compact sources in the Magellanic Clouds. II. The Small Magellanic Cloud

Jacco Th. van Loon^{1}, Joana M. Oliveira^{1}, Karl D. Gordon^{2}, G.C. Sloan^{3} and C.W. Engelbracht^{4}

1Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 3Department of Astronomy, Cornell University, Ithaca, NY 14853, USA 4Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721, USA

We present far-infrared spectra, λ=52–93 μm, obtained with the Spitzer Space Telescope in the Spectral Energy Distribution mode of its MIPS instrument, of a selection of luminous compact far-infrared sources in the Small Magellanic Cloud. These comprise nine Young Stellar Objects (YSOs), the compact H ii region N 81 and a similar object within N 84, and two red supergiants (RSGs). We use the spectra to constrain the presence and temperature of cool dust and the excitation conditions within the neutral and ionized gas, in the circumstellar environments and interfaces with the surrounding interstellar medium. We compare these results with those obtained in the LMC. The spectra of the sources in N 81 (of which we also show the ISO-LWS spectrum between 50–170 μm) and N 84 both display strong [O i] λ63-μm and [O iii] λ88-μm fine-structure line emission. We attribute these lines to strong shocks and photo-ionized gas, respectively, in a “champagne flow” scenario. The nitrogen content of these two H ii regions is very low, definitely N(N)/N(O) < 0.04 but possibly as low as N(N)/N(O) < 0.01. Overall, the oxygen lines and dust continuum are weaker in star-forming objects in the SMC than in the LMC. We attribute this to the lower metallicity of the SMC compared to that of the LMC. Whilst the dust mass differs in proportion to metallicity, the oxygen mass differs less; both observations can be reconciled with higher densities inside star-forming cloud cores in the SMC than in the LMC. The dust in the YSOs in the SMC is warmer (37–51 K) than in comparable objects in the LMC (32–44 K). We attribute this to the reduced shielding and reduced cooling at the low metallicity of the SMC. On the other hand, the efficiency of the photo-electric effect to heat the gas is found to be indistinguishable to that measured in the same manner in the LMC, ≈ 0.1–0.3%. This may result from higher cloud-core densities, or smaller grains, in the SMC. The dust associated with the two RSGs in our SMC sample is cool, and we argue that it is swept-up interstellar dust, or formed (or grew) within the bow-shock, rather than dust produced in these metal-poor RSGs themselves. Strong emission from crystalline water ice is detected in at least one YSO. The spectra constitute a valuable resource for the planning and interpretation of observations with the Herschel Space Observatory and the Stratospheric Observatory For Infrared Astronomy (SOFIA).

Accepted for publication in The Astronomical Journal Available from arXiv:1001.4487

Pre-main sequence turn-on as a chronometer for young clusters: NGC 346 as a benchmark

M. Cignoni^{1,2}, M. Tosi ^{2}, E. Sabbi^{3}, A. Nota^{3,4}, S. Degl’Innocenti^{5,6}, P.G. Prada Moroni^{5,6} and J.S. Gallagher^{7}

1Dipartimento di Astronomia, Università degli Studi di Bologna, Italy 2INAF — Osservatorio Astronomico di Bologna, Italy 3Space Telescope Science Institute, Baltimore, USA 4European Space Agency, Research and Scientific Support Department, Baltimore, USA 5Dipartimento di Fisica, Università di Pisa, Italy 6INFN — Sezione di Pisa, Italy 7Department of Astronomy, University of Wisconsin — Madison, USA

We present a novel approach to derive the age of very young star clusters, by using the Turn-On (TOn). The TOn is the point in the color–magnitude diagram (CMD) where the pre-main sequence (PMS) joins the main sequence (MS). In the MS luminosity function (LF) of the cluster, the TOn is identified as a peak followed by a dip. We propose that by combining the CMD analysis with the monitoring of the spatial distribution of MS stars it is possible to reliably identify the TOn in extragalactic star forming regions. Compared to alternative methods, this technique is complementary to the turn-off dating and avoids the systematic biases affecting the PMS phase. We describe the method and its uncertainties, and apply it to the star forming region NGC 346, which has been extensively imaged with the Hubble Space Telescope (HST). This study extends the LF approach in crowded extragalactic regions and opens the way for future studies with HST/WFC3, JWST and from the ground with adaptive optics.

Accepted for publication in ApJ Letters Available from arXiv:1001.4442

Characterizing the low-mass molecular component in the northern Small Magellanic Cloud

E. Muller^{1,2}, J. Ott^{3,4}, A. Hughes^{2,5}, J.L. Pineda^{6}, T. Wong^{7,2,8}, N. Mizuno^{1,9}, A. Kawamura^{1}, Y. Mizuno^{1}, Y. Fukui^{1}, T. Onishi^{1,10} and M. Rubio^{11}

1Department of Astrophysics, Nagoya University, Furo-cho, Chikusa-ku, Nagoya 464-8602, Japan 2Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW, 1710, Australia 3National Radio Astronomy Observatory, P.O. Box O, 1003 Lopezville Road, Socorro, NM 87801, USA 4California Institute of Technology, 1200 E. California Blvd, Caltech Astronomy, 105-24, Pasadena, CA, 91125, USA 5Centre for Supercomputing and Astrophysics, Swinburne University of Technology, Hawthorn, VIC 3122, Australia 6Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA , USA 7School of Physics, University of New South Wales, Sydney, NSW, 2052, Australia 8Department of Astronomy, University of Illinois, Urbana, IL, 61801, USA 9National Astronomical Observatory of Japan, 2-21-1 Oswa, Mitaka, Tokyo, 181-8588, Japan 10Department of Physical Science, Osaka Prefecture University, Gakuen 1-1, Sakai, Osaka 599-8531, Japan 11Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile

We present here the first results from a high-resolution survey of the ^{12}CO(J = 1−0) emission across the northern part of the poorly-enriched Small Magellanic Cloud, made with the ATNF Mopra telescope. Three molecular complexes detected in the lower resolution NANTEN survey are mapped with a beam FWHM of ∼ 42′′, to sensitivities of approximately 210 mK per 0.9 km s^{−1} channel, resolving each complex into 4–7 small clouds of masses in the range of M_{vir} ∼ 10^{3−4} M_{⊙} and with radii no larger than 16 pc. The northern SMC CO clouds follow similar empirical relationships to the southern SMC population, yet they appear relatively under-luminous for their size, suggesting that the star-forming environment in the SMC is not homogeneous. Our data also suggests that the CO cloud population has little or no extended CO envelope on scales < 30 pc, further evidence that the weak CO component in the north SMC is being disassociated by penetrating UV radiation. The new high-resolution data provide evidence for a variable correlation of the CO integrated brightness with integrated H i and μm emission; in particular CO is often, but not always, found coincident with peaks of um emission, verifying the need for matching-resolution μm and H i data for a complete assessment of the SMC H_{2} mass.

Accepted for publication in ApJ Available from arXiv:1001.4757

Conference Papers

Observations of the Large Magellanic Cloud with Fermi

J. Knödlseder^{1}, P. Jean^{1} and on behalf of the Fermi/LAT Collaboration

1Centre d’Étude Spatiale des Rayonnements, CNRS/UPS, BP 44346, F-31028 Toulouse Cedex 4, France

We report on observations of the Large Magellanic Cloud with the Fermi Gamma-Ray Space Telescope. The LMC is clearly detected with the Large Area Telescope (LAT) and for the first time the emission is spatially well resolved in γ-rays. Our observations reveal the massive star forming region 30 Doradus as a bright source of γ-ray emission in the LMC. The observations furthermore show that the γ-ray emission correlates little with the gas density of the LMC. Implications of this finding will be discussed.

Oral contribution, published in ”2009 Fermi Symposium”, Washington, D.C., Nov. 2–5 Available from arXiv:0912.4163

The luminosities of type II Cepheids and RR Lyrae variables

Michael Feast^{1}

1South African Astronomical Observatory and University of Cape Town, South Africa

Recent work on the luminosities of type II Cepheids (CephIIs) and RR Lyrae variables is reviewed. In the near infrared (JHK_{s}) the CephIIs in globular clusters show a narrow, linear, period–luminosity relation over their whole period range (∼ 1 to 100 days). The CephIIs in the general field of the LMC follow this relation for periods shorter than ∼ 20 days. At longer period (the region of the RV Tau stars), the LMC field stars have a significant scatter and in the mean are more luminous than the PL relation. The OGLEIII optical data for the LMC field variables show similar trends. Infrared colours of stars in the RV Tau period range show marked mean differences between three groupings; the Galactic field, the LMC field, and globular clusters. In the case of the Galactic field, at least, this may be strongly influenced by selection effects. In the period range ∼ 4 to 20 days (the W Vir range) there are stars lying above the PL relation which may be recognized by their light curves and are all likely to be binaries. The bright Galactic variable, κ Pav probably belongs to this group. There is evidence that CephIIs in the general field (LMC and Galaxy) have a wider range of masses than those in globular clusters. At present the CephII PL zero-point depends on the pulsation parallaxes of two stars.

Zero-points of RR Lyrae M_{V}–[Fe/H] and K_{s}–log P relations can be obtained from trigonometrical, statistical and pulsation parallaxes. These zero-points are compared with those for CephIIs and with the classical Cepheid scale using variables of these three types in the LMC. Within the uncertainties (∼ 0.1 mag) the various scales are in agreement.

Oral contribution, published in ”Variable stars, the Galactic halo and galaxy formation” (Moscow conference 2009), eds. C. Sterken, N. Samus’ & L. Szabados Available from arXiv:0912.4159

Review Paper

A universal stellar Initial Mass Function? A critical look at variations

Nate Bastian^{1}, Kevin R. Covey^{2,3} and Michael R. Meyer^{4,5}

1IoA Cambridge, UK 2Cornell, USA 3Harvard-Smithsonian CfA, USA 4ETH Zurich, Switzerland 5Steward Observatory, USA

Few topics in astronomy initiate such vigorous discussion as whether or not the initial mass function (IMF) of stars is universal, or instead sensitive to the initial conditions of star formation. The distinction is of critical importance: the IMF influences most of the observable properties of stellar populations and galaxies, and detecting variations in the IMF could provide deep insights into the process by which stars form. In this review, we take a critical look at the case for IMF variations, with a view towards whether other explanations are sufficient given the evidence. Studies of the field, local young clusters and associations, and old globular clusters suggest that the vast majority were drawn from a “universal” IMF: a power-law of Salpeter index (Γ = 1.35) above a few solar masses, and a log normal or shallower power-law (Γ ∼ 0 − 0.25) between a few tenths and a few solar masses (ignoring the effects of unresolved binaries). The shape and universality of the IMF at the stellar-substellar boundary is still under investigation and uncertainties remain large, but most observations are consistent with a IMF that declines (Γ < −0.5) well below the hydrogen burning limit. Observations of resolved stellar populations and the integrated properties of most galaxies are also consistent with a “universal IMF”, suggesting no gross variations in the IMF over much of cosmic time. There are indications of “non-standard” IMFs in specific local and extragalactic environments, which clearly warrant further study. Nonetheless, there is no clear evidence that the IMF varies strongly and systematically as a function of initial conditions after the first few generations of stars. We close with suggestions for future work that might uncover more subtle IMF variations than those that could be discerned to date.

Published in Annual Reviews of Astronomy and Astrophysics Available from arXiv:1001.2965